# Chain of Condition: Construct, Verify and Solve Conditions for Conditional Question Answering

### v1

@@ 5.1 Datasets and Evaluation Metrics
Throughout our experiments, we use two conditional question answering datasets: ConditionalQA Sun et al. (2021a) and ShARC Saeidi et al. (2018). More information about these datasets are in Appendix A.
ConditionalQA is a dataset features at long and complex documents, and has many different types of question. The document in ConditionalQA is well-structured, because it is directly crawled from websites and contains HTML tags for each paragraph. This brings the convenience for condition identification and augmentation.
We use the metrics from the original paper Sun et al. (2021a) for evaluation, which includes two sets of metrics: EM/F1 and conditional EM/F1 (abbreviated as w/conds). EM measures the exact match of predicted answer spans with gold ones, while F1 is the harmonic mean of token-level precision and recall. Conditional EM/F1 jointly measures the correctness of answer spans and the predicted conditions, providing a more comprehensive assessment of a model’s performance on the CQA task. The exact metric computation functions are in Appendix C.
ShARC is a conversational QA dataset, and the original task is to answer the question if the information in the dialog history is enough, or to generate a new question to acquire missing information. We follow the previous work Puerto et al. (2024) to isolate the QA task from the conversational setting to form a benchmark of the CQA task, resulting in a dataset that the model only needs to answer "yes", "no" or "not enough information". Additionally, we discard all irrelevant questions from the dataset for better measurement.
We evaluating model’s accuracy on the ternary classification about the answer. Since there are no human annotated conditions in the dataset, so it is not possible to further measure the accuracy of missing conditions predicted by the model, and we leave the more precise evaluation of ShARC as future work.
@@ 5.2 Baselines
We compare our approach, chain of condition, with 4 different prompting baselines in total.
$\bullet$ Code Prompting Puerto et al. (2024) is the only existing approach for prompting models for the CQA task as far as we know. This method extend the original text prompt with additional LLM-generated codes, which elicits the model’s conditional reasoning abilities for CQA tasks.
$\bullet$ Self-Ask Press et al. (2023) is a recently proposed, well-performing prompting method, and we adapt it for the CQA task. This method decompose the question by explicitly asking and answering intermediate questions until reaching the final answer and missing conditions.
Additionally, we use Zero Shot prompting and Chain of Thought prompting Wei et al. (2023) as our baselines.
@@
The supervised baselines for ConditionalQA include ETC pipeline Ainslie et al. (2020), DocHopper Sun et al. (2021b), FID Izacard and Grave (2021), SDHG Du et al. (2023), TReasoner Sun et al. (2022), and LSD Wang et al. (2023).
For ShARC, since we follow previous work to modify the dataset’s output format and discard all irrelevant instancesPuerto et al. (2024), no available supervised baselines exist. Therefore, we only compare chain of condition with other prompting baselines mentioned above.
@@ 5.3 LLM Setup
We conduct our experiments on four different large language models (LLMs) to investigate whether chain of condition performs consistently better across various settings. We use a commercial model, GPT-3.5-Turbo, and three open-source models, Llama-2-70B-chat, Llama-2-13B-chat, and Mistral-7B. Additionally, we leverage GPT-4 OpenAI (2023) for limited experiments exclusively on ConditionalQA due to cost constraints. For all models, we set the temperature to 0.0 to ensure reproducibility of the results, while using default settings for others.
The original documents in ConditionalQA can be up to 9320 tokens long, exceeding the context limitations of many LLMs, posing a challenge for all prompting methods. This could be solved by introducing a retriever to retrieve only relevant paragraphs of the document. Therefore, to address this issue and eliminate the interference from retriever performance in our experiments, we use an oracle retriever to select relevant passages for the question. We follow the methodology of previous work Puerto et al. (2024) by retaining all sections that include at least one human-annotated gold evidence and concatenating them to form the input.
And for comparison with supervised methods, we employ two approaches: (1) Using a retriever to retrieve relevant paragraphs from the document, and (2) Using a long-context version of an LLM as our backbone model. The results of these approaches will be discussed in Section 6.1.
See more setup details in Appendix B.
@@ 6.1 Main Results
The original evaluation script of ConditionalQA provides not only the overall result but also a detailed breakdown by question type in the dev set. We report the overall results here, with more detailed results available in Appendix E.
Chain of condition outperforms all prompting baselines on each dataset. It surpasses all other baselines on both datasets, establishing a new state-of-the-art. Additionally, Self-Ask also performs relatively well on ShARC, which can be attributed to the dataset’s features. The conversational format of ShARC is naturally suitable for leveraging Self-Ask, making it reasonable for it to perform better than Zero-Shot or Chain of Thought prompting.
Chain of condition outperforms all supervised baselines. With backbone models like GPT-3.5-Turbo or GPT-4, it surpasses all supervised baselines with few-shot settings. This result highlights the promising future of prompting methods for the CQA task, not only achieving better performance but also reducing the costs for fine-tuning.
Moreover, the performance of prompting GPT-3.5 with retrieved passages is much lower than prompting GPT-3.5 with gold evidence or prompting GPT-3.5-16k with the full document. This indicates that the performance of the retriever is a bottleneck that limits the effectiveness of chain of condition. It also suggests the substantial potential of it when augmented with a better retriever.
@@ 6.2 Analysis
The importance of this step lies in two aspects. First, it ensures the model identifies all possible conditions and can explicitly solve the logical expressions by external tools, improving performance on conditions. Secondly, it allows us to generate condition-aided answer, which is only feasible if all conditions are explicitly identified.
To prove the first hypothese, we conduct an ablation study on ConditionalQA because it has gold-labeled missing conditions. In this study, we prompted GPT-3.5 to first identify all conditions, then check their fulfillment, and finally indicate all unmentioned conditions implicitly through reasoning. As shown in Table 4, this ablation results in a drop of 3.2 EM score and 2.8 F1 score for answer measurement, as well as 17.4 EM score and 19.5 F1 score for joint answer and condition measurement. The performance drop is much greater when measuring both the answer and condition compared to measuring the answer alone, which indicates that removing the condition identification step leads to a much larger decrease in accuracy for conditions. Further investigation into the model’s output reveals that the average number of predicted missing conditions for conditional answers by the model increases from 1.27 to 1.67, suggesting that the model tends to judge conditions as not mentioned by the user more frequently when the condition identification step is omitted.
@@
The discussion of the second hypothesis is covered in the ablation study of the answer generation step in the following paragraphs.
@@
The removal of contextual information can hinder the model’s ability to correctly understand the meaning of a condition. To prove this, we leverage the structured document of ConditionalQA to conduct an ablation study. We remove all other paragraphs of the condition’s subsection, keeping only the original condition as input for verification. The result of conditional EM drops by 2.4 from 52.9 to 50.5, and the conditional F1 drops by 2.5 from 61.0 to 58.5 for this setting on GPT-3.5, indicating that the performance of condition prediction decreases due to reduced verification accuracy. Thus, condition augmentation would improve verification precision.
@@
In this ablation, we remove the verification results of conditions from the input of answer generation. The results are shown in Table 5. The performance drops by 3.2 EM score and 2.8 F1 score for answers, and by 2.4 EM score and 1.9 F1 score when jointly measuring answers and conditions on ConditionalQA. Additionally, the accuracy drops by 2.7 on ShARC.
Furthermore, we find that the performance drop on ConditionalQA is mostly attributed to the yes/no type questions, with a drop of 7.0 EM/F1 score and 5.3 conditional EM/F1 score. The likely reason for this phenomenon lies in the answer determination procedure: an extracted span-type answer can be found directly in the document even without verifying any condition. However, a yes/no answer must be inferred from the document along with each condition’s fulfillment. Therefore, including the conditions’ fulfillment in the prompt helps the model by reducing the need to repeatedly infer their fulfillment, allowing it to directly synthesize the information to generate the final answer.
Chain of condition performs better at indicating missing conditions.
Most questions in these CQA datasets involve identifying and solving conditions, but only a small portion of them are truly conditional. This is because, in many cases, the conditions for the answer are all satisfied by the user’s scenario, so the model only needs to give a correct judgement on whether the answer is conditional. However, when we consider only the conditional answers in the dataset, correctly addressing them becomes more challenging. This is because the model not only needs to properly generate the answer and determine it is conditional, but also precisely indicate the missing conditions.
The experimental results support this intuition. The performance of all methods on ConditionalQA [^4] ShARC does not have human-annotated conditions, so we could not experiment on it. greatly drops when considering only the conditional answers, as shown in Table 6. Besides, chain of condition consistently outperforms other prompting baselines in this setting, demonstrating its effectiveness in indicating missing conditions.
@@
We report the performance of GPT-3.5 with all prompting methods on these two groups in Table 7. Since there isn’t a metric that directly measures the quality of predicted conditions, we additionally report the F1 score of the predicted conditions. The results highlight the increased difficulty of questions involving complex conditions, and chain of condition shows much less performance degradation in this more complex group. This indicates its superior ability to handle complex conditions. We attribute this to the explicit identification of conditions and the use of a code interpreter to resolve the logical relationships between conditions.
@@ 7 Conclusion
In this work, we propose Chain of Condition, a novel prompting approach for conditional question answering. It prompts models to identify conditions with logical expressions and introduces a Python interpreter for resolution, effectively improving precision and enhancing interpretability. We conduct experiments demonstrating that chain of condition outperforms existing prompting baselines on all benchmark datasets. Additionally, when utilizing a strong backbone model, it surpasses supervised baselines. Our work analyzes the challenges associated with CQA and highlights the importance of condition identification, paving the way for future research directions.
@@
While chain of condition outperforms all baselines, it faces limitations in situations where document length exceeds the LLM’s input context capacity, primarily due to the retriever’s poor performance. This shortcoming can reduce the effectiveness of our approach in more realistic scenarios. Additionally, the multi-step prompting framework in chain of condition decomposes the original CQA task into numerous sub-tasks, leading to lower token efficiency compared to simpler prompting baselines and typically requiring more tokens to solve the entire problem. Another limitation is the scarcity of CQA datasets, which hampers further research in this area.
@@ Acknowledgments
This work is supported by NSFC (62206070).
@@ Appendix D Error Analysis
Appendix E More Detailed Results
@@
Appendix F Prompt Examples

### v2

@@ 5.1 Datasets and Evaluation Metrics
Throughout our experiments, we use two conditional question answering datasets: ConditionalQA Sun et al. (2021a) and ShARC Saeidi et al. (2018). More dataset information is in Appendix A.
ConditionalQA features long and complex documents, and has many different types of questions. Documents in ConditionalQA are usually well-structured with HTML tags for each paragraph, because they are grouped into sections and subsections describing public policies in the UK. This brings the convenience for condition identification and augmentation.
Following previous works in CQA Sun et al. (2021a), we use the original sets of metrics: EM/F1 and conditional EM/F1 (abbreviated as w/conds). EM measures the exact match of predicted answer spans with gold ones, while F1 is the harmonic mean of token-level precision and recall. Conditional EM/F1 jointly measures the correctness of answer spans and the predicted conditions, providing a more comprehensive assessment of a model’s performance on the CQA task. The calculation formulas are in Appendix C.
ShARC is a conversational QA dataset, and the original task is to answer the question if the information in the dialog history is enough, or to generate a new question to acquire missing information. We follow the previous work Puerto et al. (2024) to isolate the QA task from the conversational setting to form a benchmark of the CQA task, resulting in a dataset that the model only needs to answer "yes", "no" or "not enough information". Additionally, we discard all irrelevant questions from the dataset for better evaluation.
We use the accuracy on the answer prediction as the metric. Since there are no human annotated conditions in ShARC, so it is not possible to further measure the accuracy of predicted missing conditions. We leave more accurate evaluation for ShARC as future work.
@@ 5.2 Baselines
We compare our approach with 4 different prompting baselines in total.
$\bullet$ Code Prompting Puerto et al. (2024) is the only available prompting approach for the CQA task as far as we know. This method extends the original text prompt with additional LLM-generated code, which elicits the model’s conditional reasoning abilities for CQA tasks.
$\bullet$ Self-Ask Press et al. (2023) is a recently proposed, well-performing prompting method, which we adapt for our CQA scenario. This method decomposes the question by explicitly asking and answering intermediate questions until reaching the final answer and missing conditions.
Additionally, we use Zero Shot and Chain of Thought prompting Wei et al. (2023) as baselines.
@@
We include SDHG Du et al. (2023), TReasoner Sun et al. (2022) and LSD Wang et al. (2023) as baselines for ConditionalQA.
For ShARC, since we follow previous work to modify the dataset’s output format and discard all irrelevant instancesPuerto et al. (2024), there are no supervised baselines available. Therefore, we only compare chain of condition with other prompting baselines mentioned above.
@@ 5.3 LLM Setup
We conduct our experiments on four different LLMs to investigate whether our chain of condition performs consistently better across various settings. We use a commercial model, GPT-3.5-Turbo, and three open-source models, Llama-2-70B-chat, Llama-2-13B-chat, and Mistral-7B. Additionally, we leverage GPT-4 OpenAI (2023) for limited experiments exclusively on ConditionalQA due to cost constraints. For all models, we set the temperature to 0.0 to ensure reproducibility of the results, while using default settings for others. More details are in Appendix B.
The original documents in ConditionalQA can be up to 9,320 tokens long, exceeding the context limitations of many LLMs, posing a challenge for all prompting methods. This could be solved by introducing a retriever to retrieve only relevant paragraphs of the document. When comparing with different prompting approaches, we use an oracle retriever to select relevant passages for a given question, to eliminate the interference from the retriever performance. Following previous work Puerto et al. (2024), we retain all sections that include at least one human-annotated gold evidence and concatenating them to form the input.
And for comparison with supervised methods, we employ two approaches: (1) Using a retriever to retrieve relevant paragraphs from the document, and (2) Using a long-context version of an LLM as our backbone model.
@@ 6.1 Main Results
The original evaluation script of ConditionalQA provides a detailed breakdown by question type. We report the overall results here, with all details in Appendix F.
Chain of condition outperforms all prompting baselines on each dataset. It establishes a new state-of-the-art. Additionally, Self-Ask also performs relatively well on ShARC, which can be attributed to the dataset’s features. ShARC consists of conversations that repeatedly ask for the questioner’s information, which is very similar to the format of Self-Ask and thus naturally suitable. Therefore, it is reasonable to use Self-Ask for more fluent reasoning, leading to better performance.
Chain of condition outperforms all supervised baselines. With backbone models like GPT-3.5-Turbo or GPT-4, our framework surpasses all supervised baselines with few-shot settings. This result highlights the promising future of prompting methods for the CQA task, not only achieving better performance but also reducing the costs for fine-tuning.
Chain of condition remains effective when the input exceeds the LLM’s context limit. The result in Table 3 show that using a retriever that better explores the document structures can achieve performance comparable to supervised baselines, while long-context LLMs can surpass all supervised baselines, demonstrating the feasibility of both methods when the input goes longer. Additionally, the time and monetary costs of long-context LLMs are higher than leveraging a retriever. Given limited computational resources, we believe that developing better retrievers is a more promising direction for future research.
@@ 6.2 Analysis
We make two hypotheses about why this step is essential. The first hypothesis is that it helps the model identify all possible conditions and ensures explicitly solution for logical expressions by external tools, both improving performance on conditions. Secondly, it allows us to generate the answer with additional information about condition verification results, which is only feasible if all conditions are explicitly identified.
To prove the first hypothesis, we conduct an ablation study on ConditionalQA since it has gold-labeled missing conditions. In this study, we prompted GPT-3.5 to perform end-to-end generation, i.e., first identify all conditions, then check their fulfillment, and finally indicate all unmentioned conditions implicitly through reasoning. As shown in Table 4, this ablation results in a drop of 3.2 EM score and 2.8 F1 score for answer correctness, as well as 17.4 EM score and 19.5 F1 score for joint answer and condition correctness. The performance drop is much greater when evaluating both answers and conditions compared to evaluating answers alone, which indicates that removing the condition identification step leads to a much larger decrease in accuracy for conditions. Further investigation into the model’s output reveals that the average number of predicted missing conditions for conditional answers increases from 1.27 to 1.67, suggesting that GPT-3.5 tends to treat a condition as not mentioned by the user more frequently when the condition identification step is omitted.
@@
The second hypothesis is discussed in the ablation of the answer generation step.
@@
The removal of contextual information can hinder the model’s ability to correctly understand the meaning of a condition. To prove this, we remove all other paragraphs of the condition’s subsection, keeping only the original condition as input for verification on ConditionalQA. The result of conditional EM drops by 2.4 from 52.9 to 50.5, and the conditional F1 drops by 2.5 from 61.0 to 58.5 for this setting on GPT-3.5, indicating that the performance of condition prediction decreases due to reduced verification accuracy.
@@
EM F1 Cond EM Cond F1
Chain of Condition 64.6 73.7 52.9 61.0
w/o Interpreter 64.6 73.7 36.7 43.4
$\Delta$16.2 -17.6
Table 6: Ablation study for leveraging the Python interpreter on ConditionalQA.
@@
In this ablation, we remove the verification results of conditions from the input of answer generation. The results are shown in Table 5. The performance drops by 3.2 EM score and 2.8 F1 score for answers, and by 2.4 EM score and 1.9 F1 score when jointly evaluating answers and conditions on ConditionalQA. Additionally, the accuracy drops by 2.7 on ShARC.
Furthermore, we find that the performance drop on ConditionalQA is mostly attributed to the yes/no questions, with a drop of 7.0 EM/F1 score and 5.3 conditional EM/F1 score. The likely reason for this phenomenon lies in the answer determination procedure: a span-type answer can be extracted directly in the document even without verifying any conditions. However, a yes/no answer must be inferred from the document along with each condition’s fulfillment. Therefore, including the conditions’ fulfillment in the prompt helps the model by reducing the need to repeatedly infer their fulfillment, allowing it to directly synthesize the information to generate the final answer.
Leveraging logical expression interpreter increases accuracy for conditions.
To demonstrate the importance of utilizing external symbolic solver for logical expressions, we conduct an ablation study where the Python interpreters is replaced with GPT-3.5. Since the resolution of logical expressions does not affect the answer generation process, the EM and F1 scores for answers alone remain consistent. However, when jointly evaluating both answers and conditions, a performance drop is observed in Table 6.
Chain of condition better finds missing conditions.
Most questions in these CQA datasets involve identifying and solving conditions, but only a small portion of them are truly conditional. This is because, in many cases, the conditions for the answer are all satisfied by the user’s scenario, so the model only needs to give a correct judgement on whether the answer is conditional. However, when we consider only the conditional answers in the dataset, correctly addressing them becomes more challenging. This is because the model not only needs to properly generate the answer and determine it as conditional, but also precisely finds the missing conditions.
The performance of all methods on ConditionalQA [^4] ShARC does not have human-annotated conditions, so we could not experiment on it. greatly drops when considering only the conditional answers, as shown in Table 7. But, our chain of condition consistently outperforms other prompting baselines in this setting, demonstrating its effectiveness in finding missing conditions.
@@
We report the performance of GPT-3.5 with all prompting methods on these two groups in Table 8. Since there is not a metric that directly measures the quality of predicted conditions, we additionally report the F1 score of the predicted conditions. The results highlight the increased difficulty of questions involving complex conditions, and chain of condition shows much less performance degradation in this more challenging group. This indicates its superior ability to handle complex conditions. We attribute this to the explicit identification of conditions and the use of a code interpreter to resolve the logical relationships among conditions.
@@ 7 Conclusion
In this work, we propose Chain of condition, a novel prompting approach for conditional question answering. Our approach prompts LLMs to identify conditions, organize them into logical expressions and introduces a Python interpreter for resolution, effectively improving precision and enhancing interpretability. Experiments on all benchmark datasets show that chain of condition outperforms existing prompting baselines. Additionally, our method surpasses supervised baselines when utilizing a strong backbone model, demonstrating the promising future of prompting LLMs for CQA and paving the way for future research directions.
@@
Retriever Performance
Although our chain of condition outperforms all baselines, we find that the retriever’s performance (with a recall of around 70%) is still a limiting factor. In this paper, we simply take an commercial embedding model (text-embedding-ada-002) as our retriever, and when the document length exceeds the LLM’s input context capacity, we find this off-the-shelf retriever still overlook some necessary conditions and evidence. We believe employing a stronger retriever would achieve better performance.
Token Efficiency
The multi-step prompting framework in chain of condition decomposes the original CQA task into several sub-tasks, thus leading to lower token efficiency compared to simpler prompting baselines and typically requiring more tokens to solve the entire problem.
Exploring Intermediate Results
The generated verification results can be better exploited. For instance, if the answer type is yes/no, then it is possible to directly leverage the condition verification results to determine the answer. This approach would not only enhance accuracy but also improve token efficiency, making it a promising direction for future research. Additionally, these intermediate results could be used as training data to distill our chain of condition framework into smaller models.
@@ Acknowledgments
This work is supported in part by NSFC (62206070, 62161160339) and Beijing Science and Technology Program (Z231100007423011). We thank Chen Zhang, Mingxu Tao, Zirui Wu and Xiao Liu for their help in this work. Also, we thank Yixun Zhou and Kangcheng Luo for valuable suggestions in writing. For any correspondence, please contact Yansong Feng.
@@ Appendix C ConditionalQA Evaluation Metrics
Model GPT Llama(70b) Llama(13b) Mistral
Recall 67% 63% 52% 39%
Table 12: Performance of condition identification.
Step 1 Step 2 Step 3
GPT-3.5 (api) 2.5 2.8 1.8
Llama-2-70b (vllm) 29.9 63.2 28.7
Llama-2-13b (vllm) 6.0 6.1 3.8
Mistral 27.2 78.1 28.7
Table 13: Run-time efficiency of our method. Where step 1 stands for Condition Identification and Logical Construction, step 2 for Condition Verification, and step 3 for Expression Solution and Answer Generation.
@@ Appendix D Error Analysis
We found that most errors occur during the condition identification step. Consequently, we further investigate its performance, as shown in Table 12. The results indicate that the performance of the condition identification step remains unsatisfactory.
Appendix E Run-Time Efficiency
We report the run-time efficiency for GPT-3.5, Llama-2-70b, Llama-2-13b, and Mistral-7b averaged on two datasets in Table 13. We calculated the second per iteration (s/it) as a metric for runtime efficiency.
Appendix F More Detailed Results
@@
Appendix G Prompt Examples